OpenStudy (anonymous):

For f(x) = 2x + 8, find x when f(x) = -7.

OpenStudy (anonymous):

You again? HAHA Once again, in 2x + 8 just replace x with -7, and work it out :D

OpenStudy (shaik0124):

f(-7)=2(-7)+8 =-14+8 =-6 is answer

OpenStudy (anonymous):

im having trouble focusing and thanks once again

OpenStudy (anonymous):

Anytime :)

OpenStudy (anonymous):

just replace -7 to x then perform the operation

OpenStudy (anonymous):

^Hey, that's what I said, sort of :/ HAHAHA :)

OpenStudy (anonymous):

haha sorry

OpenStudy (anonymous):

thank you all for your help

OpenStudy (anonymous):

your welcome :)